OpenStudy (anonymous):

A rectangular lot whose perimeter is 280 ft is fenced along 3 sides. An expensive fence along the lots length cost $35 per foot. An inexpensive fencing along the 2 side widths costs only $5 per foot. the total coast of the 3 sides is 3275. what are the dimensions? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

let x = length, y = width we have this picture 

OpenStudy (anonymous):

okok

jimthompson5910 (jim_thompson5910):

the perimeter is 280, so all 3 sides add to 280 y+x+y = 280 x+2y = 280 x = 280 - 2y

jimthompson5910 (jim_thompson5910):

the expensive fence (that runs along the length x) is $35 a ft and the inexpensive fencing (that runs along the width y) is $5 a foot so the total cost is 35x + 5y + 5y = 35x + 10y

jimthompson5910 (jim_thompson5910):

we are told the total cost is $3275, so 35x + 10y = 3275 35(280-2y) + 10y = 3275 ... plug in x = 280 - 2y now solve for y

OpenStudy (anonymous):

okay hang on

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

y= 108.75?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

use that to find x

OpenStudy (anonymous):

it does? woah. you are a better explainer than my teacher!

jimthompson5910 (jim_thompson5910):

I'm glad it's all clicking now

jimthompson5910 (jim_thompson5910):

what do you get for x?

OpenStudy (anonymous):

x=62.5?

jimthompson5910 (jim_thompson5910):

perfect on both

OpenStudy (anonymous):

gracias for your help

jimthompson5910 (jim_thompson5910):

oh wait...the perimeter of the lot is 280 not the total fencing is 280 so I messed up, I'm sorry, let me fix it

OpenStudy (anonymous):

oh okay is fine

jimthompson5910 (jim_thompson5910):

the perimeter of the lot is 280, so we have 4 sides (two of them x, two of them y) that add to 280 x+y+x+y = 280 2x+2y = 280 2(x+y) = 280 x+y = 280/2 x+y = 140 y = 140 - x ------------------------------ Then we plug this into the second equation (that is still the same) and solve for x 35x + 10y = 3275 35x + 10(140-x) = 3275 35x + 1400-10x = 3275 25x + 1400 = 3275 25x = 3275 - 1400 25x = 1875 x = 1875/25 x = 75 and use this to find y y = 140 - x y = 140 - 75 y = 65 So the length is 75 and the width is 65

OpenStudy (anonymous):

I thought we only had 3 sides?

jimthompson5910 (jim_thompson5910):

I originally thought that too, but they are talking about the perimeter of the lot, not the total fencing put up

OpenStudy (anonymous):

ohhhhhhhhhhhhh

jimthompson5910 (jim_thompson5910):

yeah definitely got me too

OpenStudy (anonymous):

its a rectangle!

jimthompson5910 (jim_thompson5910):

yes it is, one side isn't fenced because it's probably against a wall already there

OpenStudy (anonymous):

the longer dimension would be 75 right

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

i only have one chance

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

perfect.it all makes sense.. Goodnight!

jimthompson5910 (jim_thompson5910):

I'm glad it does, good night

OpenStudy (anonymous):

waiiit. i have o admit I snuck an used mathway.com for a few problems could yo explain one really quick?

jimthompson5910 (jim_thompson5910):

sure and I use calculators all the time

jimthompson5910 (jim_thompson5910):

what do you need an explanation on?

OpenStudy (anonymous):

4a+7b=-8 8a+4c=56 6b+4c=-8

jimthompson5910 (jim_thompson5910):

there are a number of ways to do this, most ways lead to using fractions let's see if this method has us using the least amount of fractions

jimthompson5910 (jim_thompson5910):

solve the second equation 8a+4c=56 for c to get 8a+4c=56 4c=56 - 8a c=56/4 - 8a/4 c = 14 - 2a now plug this into the third equation 6b+4c=-8 6b+4(14 - 2a)=-8 6b + 56 - 8a = -8 -8a + 6b = -8 - 56 -8a + 6b = -64 with me so far?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

now we have the 2 equations 4a+7b=-8 -8a + 6b = -64 how do we solve this system?

OpenStudy (anonymous):

lets multiply the top equation by 2

jimthompson5910 (jim_thompson5910):

good choice that gives you 4a+7b=-8 2*(4a+7b)=2*(-8) 8a + 14b = -16

jimthompson5910 (jim_thompson5910):

so 8a + 14b = -16 -8a + 6b = -64 --------------- 0a + 20b = -80 20b = -80 ---> b = ???

jimthompson5910 (jim_thompson5910):

oh and I added those equations

OpenStudy (anonymous):

b=-4

jimthompson5910 (jim_thompson5910):

good, then use this to find a,c

jimthompson5910 (jim_thompson5910):

so go back to any equation with only a and b in them, plug in b = -4, then solve for 'a'

OpenStudy (anonymous):

4a=13

OpenStudy (anonymous):

nah thats not right

jimthompson5910 (jim_thompson5910):

a,b,c are all whole numbers so that looks off

OpenStudy (anonymous):

4a+7(-4)=-8

jimthompson5910 (jim_thompson5910):

8a + 14b = -16 8a + 14(-4) = -16 8a - 56 = -16 ... ... ... a = ???

OpenStudy (anonymous):

wait.

OpenStudy (anonymous):

oh we multiplied by 2 didnt we

jimthompson5910 (jim_thompson5910):

yes we could use any equation with a,b in it I just happened to pick that one

OpenStudy (anonymous):

a=5

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

then use this to find c

OpenStudy (anonymous):

6b+4c=-8

jimthompson5910 (jim_thompson5910):

you could use that equation or you could use c = 14 - 2a

jimthompson5910 (jim_thompson5910):

this one is much easier since c is already isolated

OpenStudy (anonymous):

c=14-10 c=4

jimthompson5910 (jim_thompson5910):

so a = 5, b = -4, c = 4

OpenStudy (anonymous):

perfect. thank you. okay this time actually goodnight lol

jimthompson5910 (jim_thompson5910):

sure thing, good night to you as well